OpenStudy (anonymous):

What is the exact value of square root of 72 over 49

OpenStudy (lgbasallote):

if you want exact use a calculator :p lol

OpenStudy (anonymous):

\[\large \frac {\sqrt{72}}{49} = \frac {\sqrt{36*2}}{49}=\frac {6 \sqrt{2}}{49} \]

OpenStudy (anonymous):

dang... that's prolly not the problem eh?

OpenStudy (lgbasallote):

yeah @dpaInc lol

OpenStudy (anonymous):

Latex sucks!@

OpenStudy (anonymous):

it is, i just have to take the square root of 49 also and i will be fine. Thanks you for helping though, i get it now

OpenStudy (anonymous):

yw... :)